# Toward Low-Latency, High-Fidelity Calibration of the LIGO Detectors with Enhanced Monitoring Tools

### v1

@@ Abstract
Accurate and reliable calibration of the Advanced LIGO detectors has enabled a plethora of gravitational-wave discoveries over the past decade, starting with the ground-breaking discovery, GW150914. Over the past ten years, the calibrated strain data from Advanced LIGO detectors has become available at a lower latency and with more reliability. In this paper, we discuss the relevant history of Advanced LIGO calibration and introduce new tools that have been developed to enable faster and more robust calibrated strain data products in the current observing run. We discuss improvements to the robustness, reliability, and accuracy of the low-latency calibration pipeline as well as the development of a new tool for monitoring the LIGO calibration in real time.
@@ 1 Introduction
The direct detection of gravitational waves (GWs) by the Advanced Laser Interferometer Gravitational-wave Observatory (LIGO) [1] in 2015 marked the beginning of a new era in observational astrophysics and fundamental physics. The first GW detection, GW150914 [2, 3], resulted from the merger of a binary black hole system, confirming a key prediction of general relativity and demonstrating the feasibility of ground-based interferometric GW detection. Since this landmark event, the field has experienced rapid progress, with hundreds of detections of binary black hole and neutron star mergers by the global GW detector network, involving Virgo [4] and KAGRA [5].
@@
In the first three observing runs, O1–O3, in total 90 confident detections were made of GWs from compact binary mergers [6]. The fourth observing run, O4, is ongoing with a nearly doubled detection rate and significant improvement in detector performance [7, 8]. After O4, major upgrades are planned in LIGO, Virgo, and KAGRA, with the aim of increasing the sensitivity, as measured by the distance to which a binary neutron star merger event would be seen at a fiducial signal-to-noise ratio, by more than a factor of two. Thus, in the upcoming fifth observing run, O5 [9], we expect an order-of-magnitude increase in the number of detections, requiring low-latency and high-quality data products to facilitate rapid astrophysical discoveries.
@@
A fundamental component of the success of GW observations is the precise calibration of interferometric data, which reconstructs the raw, digitized electrical output of the detector into an accurate and reliable measure of the dimensionless strain [10, 11, 12]. Calibration accuracy directly impacts source parameter estimation, waveform reconstruction, and tests of general relativity, making it a critical aspect of GW astronomy [13, 14]. Miscalibrations can introduce systematic biases, affecting the astrophysical interpretation of detected signals, limiting the reliability of inferred source properties, and leading to deviations from general relativity [15].
@@
During the detection of GW150914 [10] and throughout O1–O3, calibration was performed with high accuracy but at high latency, with the most precise strain data becoming available only after careful post-processing with a significant delay on the order of several months. In contrast, current calibration methodologies enable the real-time production of publishable strain data along with a well-quantified systematic error estimate on this strain data.With further sensitivity improvements in GW observatories, which allow for high signal-to-noise ratio observations, the next major challenge is the development of calibration techniques that achieve real-time calibration with an even higher accuracy, ensuring that precision science can be performed immediately upon detection. This advancement will be crucial for multi-messenger astronomy and for rapid source parameter estimation with high fidelity in future observing runs.
@@
The paper is organized as follows: section 2 provides an overview of the historical background of calibration in Advanced LIGO, section 3 discusses the challenges and calibration goals in the next observing run, section 4 describes new techniques and tools that have been implemetned to improve the accuracy and latency in LIGO calibration, section 5 discusses calibration requirements in the next generation GW detectors, and we conclude in section 6.
@@ 2 Background and Historical Overview of LIGO Calibration
In order to maintain the interferometer at its operating point for extended periods of time, the LIGO detectors depend on a multitude of control loops that actuate on various systems in response to sensor signals in real time. The primary control loop of astrophysical interest interferometrically maintains the difference in Fabry-Pérot arm cavity lengths. This differential arm, or DARM, degree of freedom is most sensitive to gravitational waves, and the control loop responsible for the DARM degree of freedom is referred to as the DARM loop. The DARM loop sensor is the photodetector that measures laser light at the output of the interferometer (anti-symmetric side of the beamsplitter) and the actuators are the quadruple pendulum actuators that move the test masses to keep the Fabry-Pérot cavities on resonance. Displacements in differential arm cavity length, whether from internal (e.g. interferometer actuators) or external (e.g. gravitational waves or noise) sources are encoded in the DARM loop error signal.
Calibration is the process of reconstructing the DARM displacement, $\Delta L_{\textrm{free}}$, from the DARM loop error signal, $d_{\textrm{err}}$. The DARM loop error signal and the DARM displacement are related by
@@
During observing runs, the detector response typically does not change significantly, but small changes in the detector response can impact the overall systematic error and uncertainty in the reconstructed $\Delta L_{\rm free}$ signal and therefore must be closely monitored. The detector response varies on time scales longer than one second, due to, for example, thermal effects, alignment fluctuations, and seismic activity. The time dependence of the response is tracked by constant measurements of the response using fiducial displacements (see section 2.1.2) at strategically selected frequencies in order to minimize any disturbances to the collection of astrophysical signals. The model of the detector response is then reconstituted, taking the tracked temporal changes into account. The final reconstructed $\Delta L_{\rm free}$ uses this reconstituted detector response in order to achieve the highest level of accuracy on the calibrated data. The effort to produce an accurate calibration comprises maintaining absolute reference standards, taking frequent measurements of the individual DARM loop components, constructing and validating a reliable model of the detector response, deploying an application of (1) to reconstruct $\Delta L_{\rm free}$, and estimating the systematic error and uncertainty associated with the detector response model.
@@ 2.1 Absolute References
The fundamental quantities that define the response of the detector are the amount of power circulating in the Fabry-Pérot cavities, the mass of the cavity mirrors (including that of the suspension that supports them), and the frequency tracking and/or timing system. The technical quantities that are critical for characterizing the control system surrounding that response are electronic, defined by functions converting current to voltage, raw to conditioned voltage, or voltage to current. These fundamental and technical quantities must be measured with high accuracy and precision in order to limit the amount of systematic error and uncertainty in the calibrated detector output.
Searches for gravitational waves are only quadratically sensitive to detector calibration systematic error or uncertainty [16, 17, 18]. Thus, collective systematic error and uncertainty in the calibrated data below 10% / 10 deg in magnitude / phase, respectively, independent of frequency and time has not limited detection. However, once detected, the process of estimating the astrophysical source properties of the gravitational wave is linear in detector calibration uncertainty and systematic error [19, 20, 21, 22, 23, 24]. As such, we now strive for a calibration systematic error and uncertainty that is less than 1% / 1 deg in magnitude / phase, as we expect signals with high signal-to-noise ratios in the upcoming observing runs. Such standards in calibration will allow for precise and accurate claims about the subtleties of the waveform’s astrophysical source properties. In order to characterize the uncertainty and systematic error at this level, we strive for our absolute reference network to have precision and accuracy less than 0.1% / 0.1 deg in magnitude / phase. This subsection highlights the work done to-date on achieving such a network.
@@ 2.1.2 Direct Force References
The amplitude of the Pcal-induced displacement is proportional to the amplitude of the modulated Pcal laser power. Thus, the accuracy of the periodic fiducial displacements is directly dependent on the accuracy of the Pcal laser power sensors. A global scheme to characterize the Pcal power sensors has been implemented to improve the accuracy of the absolute calibration in collaboration with national metrology institutes (NMIs) in the US (NIST) and Germany (PTB) and to reduce relative calibration errors between the LIGO and Virgo detectors [28].
Improvements in the Pcal measurement techniques and the automation of data analysis processes, along with NIST and PTB’s efforts to provide better calibration accuracy, has enabled the generation of fiducial displacements with only 0.3 % uncertainty for the current observing run , a significant improvement from 0.42 % in O3 [29] and 0.72% in O1 and O2 [27]. The major source of uncertainty in Pcal-induced displacement is unwanted rotation of the test mass due to the Pcal center of force position offset or interferometer beam position offset. At the LIGO Hanford site, the interferometer beams are deliberately positioned more than 20 mm from the nominal position at the center of the test mass to minimize the effect from point absorber defects [30]. However, at Livingston, the test masses were upgraded with better coatings before the start of O4. This allowed the interferometer beam to operate close to its nominal position, thus reducing the Pcal uncertainty estimate to 0.15 % from 0.3 %. It is expected that similar coating upgrades at the LIGO Hanford observatory would lead to the same level of Pcal uncertainty.
Another method to directly force the arm cavity mirrors makes use of gravity [31, 32]. These systems, referred to as the Newtonian calibrator (Ncal) or gravity field calibrators (Gcal), have been developed, tested and integrated at the Virgo [33, 34] and KAGRA [35] observatories and have shown promise of delivering displacement at sub-percent level accuracy. Ncals are a system of rotating masses that cause periodic changes in the local gravitational field applying a known gravitational force on the end test masses (ETMs). The relative standard uncertainty on the amplitude of the Ncal-induced displacements at Virgo during O4 is 0.17 %, a significant improvement over the previously reported uncertainties with the Ncal prototype [36]. LIGO constructed its own Ncal system, producing time varying forces at twice and three times the Ncal rotation frequency. During O3 it was demonstrated at LIGO Hanford that a calibrated displacement well above the detector sensitivity could be generated using this system to less than 1% relative uncertainty [37]. Using Pcal in conjunction with Ncal systems could help identify systematic errors in these systems and reduce overall displacement calibration uncertainty.
@@ 2.2 Differential Arm Length (DARM) Loop Model
Constructing a model for the DARM loop is crucial to producing an accurate and precise calibration. The DARM loop is broadly divided into two components: (1) the interferometric response of the dual-recycled Michelson interferometer with Fabry-Pérot cavities to differential displacement of the interferometer arms measured by photodetectors, known as the sensing function $C$, and (2) the feedback signals controlling the position of the test masses as part of the multi-stage quadruple pendula, known as the actuation function $A$. Together, these components form the DARM loop, or detector, response that we model as $R^{(\textrm{model})}$ in order to estimate $\Delta L_{\textrm{free}}$. The model of the detector response and estimation of the systematic error and uncertainty on measured strain have evolved in complexity since the detection of GW150914. A full description of the current models for the actuation and sensing functions can be found in B.
The mathematics to describe the signals measured by the interferometer DARM loop have been discussed in detail in [10, 11, 12]. Here, we provide a brief summary of the basics of the physical model. When the detector is operating in its lowest noise configuration, the DARM loop digital error signal, $d_{\textrm{err}}$, is linearly proportional to differential displacement of the arms, $\Delta L_{\textrm{free}}\equiv\Delta L_{x}-\Delta L_{y}$. The true response function, $R$, may be treated as a simple, single-input-single-output (SISO) control loop: the error signal is a suppressed version of the external disturbance, which can be expressed in the frequency domain as
@@
where $G$ is the feedback loop’s open loop gain transfer function, $1/(1+G)$ is the loop suppression transfer function, and $C$ is the sensing function. The differential arm displacement in units of physical length may be recast as strain on the detector, $h_{\textrm{det}}=\Delta L_{\textrm{free}}/L$, where $L$ is the average length of the arms. [^1] Note, this is not the astrophysical strain, $h_{\textrm{astro}}$, incident upon the detector; the relationship between astrophysical strain and detector strain involves a set of source- and frequency-dependent functions [39, 40, 41]. Using (2), we can solve for the more physically interesting quantity $h_{\textrm{det}}$, once again written in the frequency domain,
@@
The software used to construct all of the above components of the DARM loop model has evolved significantly since the detection of GW150914. The software implementation calculates frequency response values of DARM loop components forming $A^{\textrm{(model)}}$ and $C^{\textrm{(model)}}$. In O1-O2, the model was constructed in Matlab, as this was the native software for the most complex portion of the loop – the quadruple pendulum dynamics. Calibration products that rely on the model, however, were constructed in Python, utilizing Python packages including filter generation and and Gaussian Process Regression for estimating unknown systematic error (see section 2.3 and section 2.4 for further details).
@@ 2.4 Systematic Error and Uncertainty
which is equivalent to $h_{\textrm{det}}(f)/h_{\textrm{det}}^{\textrm{(model)}}(f)$ (from (3)). Note that $\eta_{R}(f)$ is a multiplicative factor to scale $h_{\textrm{det}}^{\textrm{(model)}}(f)$ to $h_{\textrm{det}}(f)$, where $h_{\textrm{det}}(f)$ is the exact detector strain that calibration is striving to reconstruct. A preliminary version of this equation appears as equation 13 in [10], and is later refined in equation 5 of [11], and solidified in equations 9 – 11 of [12].
@@
At the time of GW150914, the systematic error was modeled in a frequentist framework in the fashion of the iLIGO error model [46]. The response function was analytically broken down into partial derivatives of its components to form a “propagation of errors” assuming independence of all measurements or statements that informed the production of $A^{\textrm{(model)}}$ and $C^{\textrm{(model)}}$; see equations 7 and 8 of [11]. It became apparent, however, that the detector response was more complex than anticipated and our measurements of seemingly independent parameters showed correlation. Inspired by astrophysical parameter estimation techniques which employ numerically evaluated Bayesian statistics to characterize uncertainty, we began to create an analysis framework for modeling systematic error that appropriately handled many cross-correlated parameters with physics-agnostic algorithms necessary for the accurate representation of the error. This framework, described in [12], has been used throughout O1-O3 to model systematic error on the reconstructed detector strain. During O4, this framework has evolved using newly available measurements, described in section 4.
@@
Here we summarize the framework used in O1-O3 as it still remains within the O4 framework. Using (6), we observe that $R$ is composed of $C$ and $A$, and conclude that $\eta_{R}$ is also composed of $\eta_{C}$ and $\eta_{A}$. Section 3.2 of [12] discusses how the Pcals may be used to directly measure $A$ and $C$ for any given time, and thus build a measurement of $\eta_{C}$ and $\eta_{A}$ at the limit of Pcal uncertainty. We find that these terms are dominated by the systematic error between the models and measurements of $C$ and $A$. Section 4.3 of [12] describes the use of Gaussian Process Regression (GPR) methods to fit the measured systematic error in $\eta_{C}$ and $\eta_{A}$. Crucially, the GPR provides a posterior distribution of the fits of $\eta_{C}$ and $\eta_{A}$ that we sample numerically as part of the estimate of $\eta_{R}$. This results in a posterior distribution of $\eta_{R}$ at each selected short interval of time. [^2] This is typically done during a 2-minute interval at an hourly cadence while the detector is in a mode suitable for observation, but the process can be run at specific times, if desired. Although we implicitly assume that successive measurements of $\eta_{C}$ or $\eta_{A}$ are drawn from the same parent distribution, in reality the detectors evolve in time, and our models for $C$ and $A$ may not fully reflect the evolution. Nevertheless, kernel hyper parameter values in the GPR are chosen such that that the measurement results are contained within the 1-sigma percentile of the distribution of curves instantiated from the GPR kernel.
@@
In [10], we published the model of systematic error in the reconstructed strain data at the time of GW150914. By the end of O1, the number of transient GW detections were still low enough to again publish the specific systematic error estimates for the detector strain at the time of each detection, see Appendix A of [11]. During O2, although astrophysical parameter searches still used event-by-event systematic error estimates, it became necessary to quantify the systematic error estimate over several entire epochs of observation time / detector configurations  [11]. The procedure remained in place for O3 [12, 47], with seven epochs for the Hanford detector and four epochs for the Livingston detector. Epochs are defined by major detector changes throughout the run. Experience now dictates that detectors will have many such epochs throughout observing runs. The measurement sets that define $\eta_{R}$ within those epochs – the measured systematic error in $\eta_{C}$ and $\eta_{A}$ – may have entirely different response after an epoch boundary. As such, the measurement collection is “reset” every boundary and we do not expect $\eta_{R}$ to improve linearly with time.
@@
Delivering the model of systematic error has historically taken several weeks to months, due to the complexity of combining measurements, accounting for detector changes, and addressing larger known systematics. During O1-O3, the strain data reconstructed in low-latency was not accompanied by a systematic error estimate. Instead, the systematic error estimate was determined from a revised version of the strain data generated offline, with latency on the order of months. During O4, we have addressed many of the challenges so that systematic error estimates are now provided alongside the low-latency calibrated strain data.

### v2

@@ Abstract
Accurate and reliable calibration of the Advanced LIGO detectors has enabled a plethora of gravitational-wave discoveries in the detectors’ first decade of operation, starting with the ground-breaking discovery, GW150914. In the first decade of operation, the calibrated strain data from Advanced LIGO detectors has become available at a lower latency and with more reliability. In this paper, we discuss the relevant history of Advanced LIGO calibration and introduce new tools that have been developed to enable faster and more robust calibrated strain data products in the fourth observing run (O4). We discuss improvements to the robustness, reliability, and accuracy of the low-latency calibration pipeline as well as the development of a new tool for monitoring the LIGO detector calibration in real time.
@@ 1 Introduction
The direct detection of gravitational waves (GWs) by the Advanced Laser Interferometer Gravitational-wave Observatory (LIGO) [1] detector in 2015 marked the beginning of a new era in observational astrophysics and fundamental physics. The first GW detection, GW150914 [2, 3], resulted from the merger of a binary black hole system, confirming a key prediction of general relativity and demonstrating the feasibility of ground-based interferometric GW detection. Since this landmark event, the field has experienced rapid progress, with hundreds of detections of binary black hole and neutron star mergers by the global GW detector network of LIGO, Virgo [4], and KAGRA [5].
@@
In the first three observing runs, O1–O3, in total 90 confident detections were made of GWs from compact binary mergers [6]. The fourth observing run, O4, is ongoing at the writing of this paper with a nearly doubled detection rate and significant improvement in detector performance [7, 8]. After O4, major upgrades are planned in LIGO, Virgo, and KAGRA, with the aim of increasing the sensitivity, as measured by the distance to which a binary neutron star merger event would be seen at a fiducial signal-to-noise ratio, by more than a factor of two. Thus, in the upcoming fifth observing run, O5 [9], we expect an order-of-magnitude increase in the number of detections, requiring low-latency and high-quality data products to facilitate rapid astrophysical discoveries.
@@
A fundamental component of the success of GW observations is the precise calibration of interferometric data, which reconstructs the raw, digitized electrical output of the detector into an accurate and reliable measure of the dimensionless strain [10, 11, 12]. Calibration accuracy directly impacts source parameter estimation, waveform reconstruction, and tests of general relativity, making it a critical aspect of GW astronomy [13, 14]. Miscalibrations can introduce systematic biases, affecting the astrophysical interpretation of detected signals, limiting the reliability of inferred source properties, and leading to false deviations from general relativity [15].
@@
At the time of the detection of GW150914 [10] and throughout O1–O3, calibration was performed with high accuracy but at high latency, with the most precise strain data becoming available only after careful post-processing with a significant delay on the order of several months. In contrast, calibration methodologies implemented in O4 enable the real-time production of publishable strain data along with a well-quantified systematic error estimate on this strain data.Figure 1 summarizes the evolution of various features in the LIGO calibration process from O1 through O4, which are discussed in more detail in section 4. With further sensitivity improvements in GW observatories, which allow for high signal-to-noise ratio observations, the next major challenge is the development of calibration techniques that achieve real-time calibration with an even higher accuracy, ensuring that precision science can be performed immediately upon detection. This advancement will be crucial for multi-messenger astronomy and for rapid source parameter estimation with high fidelity in future observing runs.
@@
Although LIGO, Virgo, and KAGRA share similar core calibration schemes [16, 5], each has developed distinct calibration pipelines and adopts different combinations of absolute references [17, 18, 19, 20]. This paper focuses on the methods and developments specific to LIGO. The paper is organized as follows: section 2 provides a robust overview of the historical background of calibration in Advanced LIGO over the decade following the discovery of GW150914, section 3 discusses the expected challenges and calibration goals in the next observing run (O5), section 4 describes new techniques and tools that have been implemented to improve the accuracy and latency of calibrated LIGO data, section 5 discusses calibration requirements in the next generation GW detectors, and we conclude in section 6.
Figure 1: Schematic visualizing the evolution of several aspects of LIGO calibration from O1 through O4. This graphic focuses on five specific components of the LIGO calibration procedure which are discussed primarily in section 4.
@@ 2 Background and Historical Overview of LIGO Calibration
In order to maintain the interferometer at its operating point for extended periods of time, the LIGO detectors depend on a multitude of control loops that actuate on various systems in response to sensor signals in real time. The primary control loop of astrophysical interest interferometrically maintains the difference in Fabry-Perot arm cavity lengths. This differential arm, or DARM, degree of freedom is most sensitive to gravitational waves, and the control loop responsible for the DARM degree of freedom is referred to as the DARM loop. The DARM loop sensor is the photodetector that measures laser light at the output of the interferometer (anti-symmetric side of the beamsplitter) and the actuators are the quadruple pendulum actuators that move the test masses to keep the Fabry-Perot cavities on resonance. Displacements in differential arm cavity length, whether from internal (e.g. interferometer actuators) or external (e.g. gravitational waves or noise) sources are encoded in the DARM loop error signal.
Calibration is the process of reconstructing the DARM displacement, $\Delta L_{\textrm{free}}$, from the DARM loop error signal, $d_{\textrm{err}}$. In the frequency domain, the DARM loop error signal and the DARM displacement are related by
@@
During observing runs, the detector response typically does not change significantly, but small changes can impact the overall systematic error and uncertainty in the reconstructed $\Delta L_{\rm free}$ signal and therefore must be closely monitored. The detector response varies on time scales longer than one second, due to, for example, thermal effects, alignment fluctuations, and seismic activity. The time-dependence of the response is tracked by constant measurements of the response using fiducial displacements (see section 2.1.2) at strategically selected frequencies in order to minimize any disturbances to the collection of astrophysical signals. The model of the detector response is then reconstituted, taking the tracked temporal changes into account. The final reconstructed $\Delta L_{\rm free}$ uses this reconstituted detector response in order to achieve the highest level of accuracy on the calibrated data. The effort to produce an accurate calibration comprises maintaining absolute reference standards, taking frequent measurements of the individual DARM loop components, constructing and validating a reliable model of the detector response, deploying an application of (1) to reconstruct $\Delta L_{\rm free}$, and estimating the systematic error and uncertainty associated with the detector response model.
@@ 2.1 Absolute References
The fundamental quantities that define the response of the detector are the amount of power circulating in the Fabry-Perot cavities, the mass of the cavity mirrors (including that of the suspension that supports them), and the frequency tracking and/or timing system. The technical quantities that are critical for characterizing the control system surrounding that response are electronic, defined by functions converting current to voltage, raw to conditioned voltage, or voltage to current. These fundamental and technical quantities must be measured with high accuracy and precision in order to limit the amount of systematic error and uncertainty in the calibrated detector output.
Searches for gravitational waves are only quadratically sensitive to detector calibration systematic error or uncertainty [21, 22, 23]. Thus, collective systematic error and uncertainty in the calibrated data below 10% in magnitude and 10 deg in phase, independent of frequency and time has not limited detection. However, once detected, the process of estimating the astrophysical source properties of the gravitational wave is linear in detector calibration uncertainty and systematic error [24, 25, 26, 27, 28, 29]. Additionally, calibration uncertainty has been shown to always reduce the sensitivity of modeled searches for gravitational waves [30]. As such, we now strive for a calibration systematic error and uncertainty that is less than 1% in magnitude and 1 deg in phase, as we expect signals with high signal-to-noise ratios in the upcoming observing runs. Such standards in calibration will allow for precise and accurate claims about the subtleties of astrophysical source properties derived from gravitational waveforms. In order to characterize the uncertainty and systematic error at this level, we strive for our absolute reference network to have precision and accuracy less than 0.1% in magnitude and 0.1 deg in phase. This subsection highlights the work done to-date on achieving such a network.
@@ 2.1.2 Direct Force References
The amplitude of the Pcal-induced displacement is proportional to the amplitude of the modulated Pcal laser power. Thus, the accuracy of the periodic fiducial displacements is directly dependent on the accuracy of the Pcal laser power sensors. A global scheme to characterize the Pcal power sensors has been implemented to improve the accuracy of the absolute calibration in collaboration with national metrology institutes in the US (NIST) and Germany (PTB) and to reduce relative calibration errors between the LIGO and Virgo detectors [35].
Improvements in the Pcal measurement techniques and the automation of data analysis processes, along with NIST and PTB’s efforts to provide better calibration accuracy, has enabled the generation of fiducial displacements with only 0.3 % uncertainty for the current observing run, a significant improvement from 0.42 % in O3 [36] and 0.72% in O1 and O2 [34]. The major source of uncertainty in Pcal-induced displacement is unwanted rotation of the test mass due to the Pcal center of force position offset or an interferometer beam position offset. At the LIGO Hanford site, the interferometer beams are deliberately positioned more than 20 mm away from the nominal position at the center of the test mass to minimize the effect from point absorber defects [37]. However, at Livingston, the test mass coatings were improved to remove point absorber defects before the start of O4. This allowed the interferometer beam to operate closer to its nominal position, thus reducing the Pcal uncertainty estimate from 0.3 % to 0.15 %. It is expected that similar upgrades to reduce point absorbers on the coatings at the LIGO Hanford observatory would lead to the same level of Pcal uncertainty.
Another method to directly apply a force to the arm cavity mirrors makes use of gravity [38, 39]. These systems, referred to as the Newtonian calibrator (Ncal) or gravity field calibrators (Gcal), have been developed, tested and integrated at the Virgo [18, 19] and KAGRA [20] observatories and have shown promise of delivering displacement at sub-percent level accuracy. Ncals are a system of rotating masses that cause periodic changes in the local gravitational field applying a known gravitational force on the end test masses (ETMs). The relative standard uncertainty on the amplitude of the Ncal-induced displacements at Virgo during O4 is 0.17 %, a significant improvement over the previously reported uncertainties with the Ncal prototype [17]. LIGO has constructed its own Ncal system, producing time varying forces at twice and three times the Ncal rotation frequency. During O3 it was demonstrated at LIGO Hanford that a calibrated displacement well above the detector sensitivity could be generated using this system to less than 1% relative uncertainty [40]. Using Pcal in conjunction with Ncal systems could help identify systematic errors in these systems and reduce overall displacement calibration uncertainty.
@@ 2.2 Differential Arm Length (DARM) Loop Model
Constructing a model for the DARM loop is crucial to producing an accurate and precise calibration. The DARM loop is broadly divided into two components: (1) the interferometric response of the dual-recycled Michelson interferometer with Fabry-Perot cavities to differential displacement of the interferometer arms measured by photodetectors, known as the sensing function $C$, and (2) the feedback signals controlling the position of the test masses as part of the multi-stage quadruple pendula, known as the actuation function $A$. Together, these components form the DARM loop, or detector, response that we model as $R^{(\textrm{model})}$ in order to estimate $\Delta L_{\textrm{free}}$. The model of the detector response and estimation of the systematic error and uncertainty on measured strain have evolved in complexity since the detection of GW150914. A full description of the current models for the actuation and sensing functions can be found in B.
The mathematics to describe the signals measured by the interferometer DARM loop have been discussed in detail in [10, 11, 12]. Here, we provide a brief summary of the basics of the physical model. When the detector is operating in its lowest-noise configuration, the DARM loop digital error signal, $d_{\textrm{err}}$, is linearly proportional to differential displacement of the arms, $\Delta L_{\textrm{free}}\equiv\Delta L_{x}-\Delta L_{y}$. The true response function, $R$, may be treated as a simple, single-input-single-output (SISO) control loop: the error signal is a suppressed version of the external disturbance, which can be expressed in the frequency domain as
@@
where $G$ is the feedback loop’s open loop gain transfer function, $1/(1+G)$ is the loop suppression transfer function, and $C$ is the sensing function. The differential arm displacement in units of physical length may be recast as strain on the detector, $h_{\textrm{det}}=\Delta L_{\textrm{free}}/L$, where $L$ is the average length of the arms [^1] Note, this is not the astrophysical strain, $h_{\textrm{astro}}$, incident upon the detector; the relationship between astrophysical strain and detector strain involves a set of source- and frequency-dependent functions [42, 43, 44]. . Using (2), we can solve for the more physically interesting quantity $h_{\textrm{det}}$, once again written in the frequency domain,
@@
The software used to construct all of the above components of the DARM loop model has evolved significantly since the detection of GW150914. The software implementation calculates frequency response values of DARM loop components forming $A^{\textrm{(model)}}$ and $C^{\textrm{(model)}}$. In O1-O2, the model was constructed in Matlab, as this was the native software for the most complex portion of the loop – the quadruple pendulum dynamics. Calibration products that rely on the model, however, were constructed in Python, utilizing Python packages including filter generation and Gaussian Process Regression for estimating unknown systematic error (see section 2.3 and section 2.4 for further details).
@@ 2.4 Systematic Error and Uncertainty
which is equivalent to $h_{\textrm{det}}(f)/h_{\textrm{det}}^{\textrm{(model)}}(f)$ (from (3)). Note that $\eta_{R}(f)$ is a multiplicative factor to scale $h_{\textrm{det}}^{\textrm{(model)}}(f)$ to $h_{\textrm{det}}(f)$, where $h_{\textrm{det}}(f)$ is the exact detector strain that calibration is striving to reconstruct. A preliminary version of this equation appears as equation (13) in [10], and is later refined in equation (5) of [11], and solidified in equations (9) – (11) of [12].
@@
At the time of GW150914, the systematic error was modeled in a frequentist framework in the fashion of the iLIGO error model [49]. The response function was analytically broken down into partial derivatives of its components to form a “propagation of errors” assuming independence of all measurements or statements that informed the production of $A^{\textrm{(model)}}$ and $C^{\textrm{(model)}}$; see equations (7) and (8) of [11]. It became apparent, however, that the detector response was more complex than anticipated and our measurements of seemingly independent parameters showed correlation. Inspired by astrophysical parameter estimation techniques which employ numerically evaluated Bayesian statistics to characterize uncertainty, we began to create an analysis framework for modeling systematic error that appropriately handled many cross-correlated parameters with physics-agnostic algorithms necessary for the accurate representation of the error. This framework, described in [12], has been used throughout O1-O3 to model systematic error on the reconstructed detector strain. During O4, this framework has evolved using newly available measurements, described in section 4.
@@
Here we summarize the framework used in O1-O3 as it still remains within the O4 framework. Using (6), we observe that $R$ is composed of $C$ and $A$, and conclude that $\eta_{R}$ is also composed of $\eta_{C}$ and $\eta_{A}$. Section 3.2 of [12] discusses how the Pcals may be used to directly measure $A$ and $C$ for any given time, and thus build a measurement of $\eta_{C}$ and $\eta_{A}$ at the limit of Pcal uncertainty. We find that these terms are dominated by the systematic error between the models and measurements of $C$ and $A$. Section 4.3 of [12] describes the use of Gaussian Process Regression (GPR) methods [50] to fit the measured systematic error in $\eta_{C}$ and $\eta_{A}$. Crucially, the GPR provides a posterior distribution of the fits of $\eta_{C}$ and $\eta_{A}$ that we sample numerically as part of the estimate of $\eta_{R}$. This results in a posterior distribution of $\eta_{R}$ at each selected short interval of time [^2] This is typically done during a 2-minute interval at an hourly cadence while the detector is in a mode suitable for observation, but the process can be run at specific times, if desired.. Although we implicitly assume that successive measurements of $\eta_{C}$ or $\eta_{A}$ are drawn from the same parent distribution, in reality the detectors evolve in time, and our models for $C$ and $A$ may not fully reflect the evolution. Nevertheless, kernel hyperparameter values, which encode information about the shape and behavior of the covariance function used in the GPR, are chosen such that the measurement results are contained within the 1-$\sigma$ percentile of the distribution of curves instantiated from the GPR kernel.
@@
In [10], we published the model of systematic error in the reconstructed strain data at the time of GW150914. By the end of O1, the number of transient GW detections were still low enough to again publish the specific systematic error estimates for the detector strain at the time of each detection, see Appendix A of [11]. During O2, although astrophysical parameter searches still used event-by-event systematic error estimates, it became necessary to quantify the systematic error estimate over several entire epochs of observation time [11]. The procedure remained in place for O3 [12, 51], with seven epochs for the Hanford detector and four epochs for the Livingston detector. Epochs are defined by major detector changes throughout the run. The measurement sets that define $\eta_{R}$ within those epochs – the measured systematic error in $\eta_{C}$ and $\eta_{A}$ – may have entirely different response after an epoch boundary. As such, the measurement collection is “reset” at every boundary and we do not expect $\eta_{R}$ to improve linearly with time.
@@
Delivering the model of systematic error has historically taken several weeks to months, due to the complexity of combining measurements, accounting for detector changes, and addressing larger known systematics. During O1-O3, the strain data reconstructed in low-latency was not accompanied by a systematic error estimate (see figure 1). Instead, the systematic error estimate was determined from a revised version of the strain data generated offline, with latency on the order of months. During O4, we have addressed many of the challenges so that systematic error estimates are now provided alongside the low-latency calibrated strain data.